Shift report 20010923 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Sep 23 2001 - 16:41:53 EDT

  • Next message: Apache: "Shift report 20010924 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Sep 23 16:41:53 2001, Shift: 08:00-16:00
    Supervisor: JH Lee, with: 
    Log book#9 pages 30 - 31
    
    Shift Summary:
    ------------------------------------------------------------
    1. Cogged Beam until 11:30 
    
    2. Spectrometer Settings:  MRS@90 4A
                               FS@8 1/5B
       
    3. Data:   
       Drift Chamber (T3,T4,T5) HV Scan runs (#4876-#4879)
       with 40V step from the default HV settings.
       run4876:  80V less than default HV
       run4877: 120V less
       run4878: 160V less
       run4879:  40V less
    
      (run4875:  50V less: 
       see the shift report from the previous shift)      
     
    3. FS at 8 degree program finished
       FS moved to 3 degree (at 1/3 field)
    
    4. No new detector problem
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Sep 23 2001 - 16:42:33 EDT